# On K-peak solutions for the Yamabe equation on product manifolds

### 09 Mar 2026 v1

@@ 4. Asymptotic expansion of the energy functional
(71) $c_{1}:=\frac{1}{6}\int_{{\mathbb{R}}^{n}}\left(\frac{U^{\prime}(|z|)}{|z|}\right)^{2}z_{1}^{4}dz$
(72) $c_{2}:=\int_{{\mathbb{R}}^{n}}U^{2}z_{1}^{2}dz.$
(73) $c_{3}:=\int_{{\mathbb{R}}^{n}}z_{1}^{2}\psi_{1}(|z|)\frac{U^{\prime}(|z|)}{|z|}dz$
(74) $c_{4}:=\int_{{\mathbb{R}}^{n}}z_{1}^{2}\psi_{2}(|z|)\frac{U^{\prime}(|z|)}{|z|}dz$
(75) $c_{5}:=\frac{{\bf c}}{6n}\int_{{\mathbb{R}}^{n}}\left(\frac{1}{2}(U^{\prime}(z))^{2}-\frac{1}{2-p}\frac{U(z)U^{\prime}(z)}{|z|}\right)z_{1}^{2}dz-\frac{{\bf c}}{6n}\int_{{\mathbb{R}}^{n}}U(z)\psi_{2}(|z|)dz$
(76) $c_{6}:=8c_{1}-\frac{40c_{3}}{3}$
(77) $c_{7}:=-\frac{c_{4}}{18n}-c_{5}-\frac{c_{2}\mathbf{c}}{12}+\frac{c_{1}}{24(n+2)}$
@@
$\frac{1}{\epsilon^{n}}\frac{1}{2}\int_{M}s_{g}W_{\epsilon_{i},\xi_{i}}^{2}d\mu_{g}=\frac{1}{2}\int_{B(0,r/\epsilon)}s_{g}(\exp_{\xi}(\epsilon z))(U(z)\chi_{r/\epsilon}(\epsilon z))^{2}\sqrt{det(g_{\xi}(\epsilon z))}dz$
@@
Also
$\int_{{\mathbb{R}}^{n}}U^{2}z_{j}dz=0$
This implies
@@
$\sum_{i=1}^{K}\ \epsilon^{5}\frac{1}{2}\frac{1}{\epsilon^{n}}\int_{M}\mathbf{c}s_{g}V_{\epsilon,\xi_{i}}^{2}d\mu_{g}+o(\epsilon^{4})$
@@
Note that by construction of $V$, eq. (21), and eq. (6) for $U$, the term
@@
$\epsilon^{2}\frac{1}{\epsilon^{n}}\int_{M}\left(-\Delta_{g}W_{\epsilon,\xi}+(1+{\bf c}s_{g}\epsilon^{2})W_{\epsilon,\xi}-|W_{\epsilon,\xi}|^{p-1}\right)V_{\epsilon,\xi}\ d\mu_{g}=-\epsilon^{4}\int_{{\mathbb{R}}^{n}}L_{0}VVdz+o(\epsilon^{4})$
Meanwhile, using the fact that $\Delta_{g}V-\Delta_{{\mathbb{R}}^{n}}V=o(\epsilon^{2})$ and rewriting (88) on ${\mathbb{R}}^{n}$:
@@
Hence, (88) turns into
@@
$\int_{{\mathbb{R}}^{n}}L_{0}V_{1}V_{1}dz$ was computed in [8] (eqs. (68) to (71) in [8]):
@@
with $c_{3}$ and $c_{4}$ dimensional constants, given by
(92) $c_{3}:=\int_{{\mathbb{R}}^{n}}\psi_{1}(|z|)\frac{U^{\prime}(|z|)}{|z|}z_{1}^{2}dz{\text{ \ \ and \ \ }}c_{4}:=\int_{{\mathbb{R}}^{n}}\psi_{2}(|z|)\frac{U^{\prime}(|z|)}{|z|}z_{1}^{2}dz$
where $\psi_{1}(|z|)$ and $\psi_{2}(|z|)$ are the functions given in (27). We now compute the rest of the terms in (91). Note first that for a radial function $\varphi(|z|)$ on ${\mathbb{R}}^{n}$,
$\int_{{\mathbb{R}}^{n}}\varphi(|z|)z_{k}z_{l}dz=\delta_{kl}\int_{{\mathbb{R}}^{n}}\varphi(|z|)z_{1}^{2}dz$
By direct computation, by eq. (28) for $L_{0}V_{1}$ and eq. (31) for $V_{2}$,
@@
$\frac{{\bf c}}{3n}s_{g}^{2}\int_{{\mathbb{R}}^{n}}\left(\frac{1}{2}(U^{\prime}(z))^{2}-\frac{1}{2-p}\frac{U(z)U^{\prime}(z)}{|z|}\right)z_{1}^{2}dz$
since $\left(\frac{1}{2}(U^{\prime}(z))^{2}-\frac{1}{2-p}\frac{U(z)U^{\prime}(z)}{|z|}\right)$ is a radial function. On the other hand, by eq. (26) for $L_{0}V_{2}$ and eq. (27) for $V_{1}$,
$\int_{{\mathbb{R}}^{n}}L_{0}V_{2}V_{1}dz=-\frac{1}{3}{\bf c}s_{g}\sum_{k\neq l}R_{kl}\int_{{\mathbb{R}}^{n}}U(z)\psi_{1}(|z|)z_{k}z_{l}dz-\frac{1}{3n}{\bf c}s_{g}^{2}\int_{{\mathbb{R}}^{n}}U(z)\psi_{2}(|z|)dz$
$\frac{1}{3n}{\bf c}s_{g}^{2}\int_{{\mathbb{R}}^{n}}U(z)\psi_{2}(|z|)dz$
since $\int_{{\mathbb{R}}^{n}}U(z)\psi_{1}(|z|)z_{k}z_{l}dz=0$ for $k\neq l$, being $U(z)\psi_{1}(|z|)$ a radial function. And by eq. (26) for $L_{0}V_{2}$ and eq. (31) for $V_{2}$,
@@
We then get, from these equationsn and (90) together with (91),
@@
(94) $c_{5}:=\frac{{\bf c}}{6n}\int_{{\mathbb{R}}^{n}}\left(\frac{1}{2}(U^{\prime}(z))^{2}-\frac{1}{2-p}\frac{U(z)U^{\prime}(z)}{|z|}\right)z_{1}^{2}dz-\frac{{\bf c}}{6n}\int_{{\mathbb{R}}^{n}}U(z)\psi_{2}(|z|)dz$
@@
with $\psi_{2}(|z|)$ the radial function from eq. (31). This implies
[…]
On the other hand, $K_{3}$, $K_{4}$, $K_{5}$ and $K_{6}$ are $o(\epsilon^{4})$ by Lemmas 6.3, 6.4, 6.6 and 6.7, respectively. It follows that
(97) $J_{\epsilon}\left(\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}+\epsilon^{2}\sum_{i=1}^{K}V_{\epsilon,\xi_{i}}\right)=K\alpha+\epsilon^{2}\frac{1}{2}\beta\sum_{i=1}^{K}s_{g}(\xi_{i})$
[…]
$\frac{1}{2}\sum_{i\neq j}^{K}\gamma_{ij}U\left(\frac{\exp_{\xi_{i}}^{-1}(\exp_{\xi_{j}}(\epsilon z))}{\epsilon}\right)+o(\epsilon^{4})$
We conclude,
@@
where, $c_{6}=8c_{1}-\frac{40c_{3}}{3}$, $c_{7}=-\frac{c_{4}}{18n}-c_{5}-\frac{c_{2}\mathbf{c}}{12}+\frac{c_{1}}{24(n+2)}$ and $c_{8}=18c_{1}+30{\bf c}c_{2}(n+2)$.
We now prove Proposition 3.2.
@@
We write $F(u)=\frac{1}{p}(u^{+})^{p}$ and $f(u)=(u^{+})^{p-1}$ (so that $F^{\prime}(u)=f(u)$). First, since $\phi_{\epsilon,\bar{\xi}}\in K^{\perp}_{\epsilon,\bar{\xi}}$,
@@ 5. The Finite dimensional reduction
We follow [11, Proposition 3.1], [18, Lemma 6.1]. We proceed by contradiction. Suppose there is a sequence $\epsilon_{j}\rightarrow 0$, $\{\epsilon_{j}\}_{j\in{\mathbb{N}}}$ together with $\{\bar{\xi}_{j}\}_{j\in{\mathbb{N}}}\subset M^{K}$, such that the hypothesis of the Lemma, eq. (102) is satisfied and at the same time there is some $\{\phi_{j}\}_{j\in{\mathbb{N}}}\subset K^{\perp}_{\epsilon,\bar{\xi}}$, with

### 19 Mar 2026 v2

@@ 4. Asymptotic expansion of the energy functional
(69) $c_{1}:=\frac{1}{6}\int_{{\mathbb{R}}^{n}}\left(\frac{U^{\prime}(z)}{|z|}\right)^{2}z_{1}^{4}dz\textnormal{, \ \ \ }c_{2}:=\int_{{\mathbb{R}}^{n}}U^{2}(z)z_{1}^{2}dz,$
(70) $c_{3}:=\frac{1}{54}\int_{{\mathbb{R}}^{n}}\psi(z)\frac{U^{\prime}(z)}{|z|}z_{1}^{4}dz,\textnormal{\ \ \ \ \ }c_{4}:=-\frac{{\bf c}}{6}\int_{{\mathbb{R}}^{n}}U(z)\psi(z)z_{1}^{2}dz$
(71) $c_{5}:=\frac{{\bf c}}{6}\int_{{\mathbb{R}}^{n}}\left(\frac{1}{2}(U^{\prime}(z))^{2}-\frac{1}{2-p}\frac{U(z)U^{\prime}(z)}{|z|}\right)z_{1}^{2}dz$
(72) $c_{6}:=8c_{1}-120(n+2)c_{3}\textnormal{,\ \ \ \ \ \ }c_{7}:=-c_{3}-c_{4}-c_{5}-\frac{c_{2}\mathbf{c}}{12}+\frac{c_{1}}{24(n+2)}$
@@
$\frac{1}{\epsilon^{n}}\frac{1}{2}\int_{M}s_{g}W_{\epsilon,\xi_{i}}^{2}d\mu_{g}=\frac{1}{2}\int_{B(0,r/\epsilon)}s_{g}(\exp_{\xi}(\epsilon z))(U(z)\chi_{r/\epsilon}(\epsilon z))^{2}\sqrt{det(g_{\xi}(\epsilon z))}dz$
@@
Also, for any $j\in\{1,2,\dots,n\}$, $\int_{{\mathbb{R}}^{n}}U^{2}z_{j}dz=0$, since $U$ is radial. This implies
@@
$\sum_{i=1}^{K}\ \epsilon^{6}\frac{1}{2}\frac{1}{\epsilon^{n}}\int_{M}\mathbf{c}s_{g}V_{\epsilon,\xi_{i}}^{2}d\mu_{g}+o(\epsilon^{4})$
@@
Note that by construction of $V$, eq. (21), and using eq. (6) for $U$, the term
@@
(85) $\epsilon^{2}\frac{1}{\epsilon^{n}}\int_{M}\left(-\Delta_{g}W_{\epsilon,\xi}+(1+{\bf c}s_{g}\epsilon^{2})W_{\epsilon,\xi}-|W_{\epsilon,\xi}|^{p-1}\right)V_{\epsilon,\xi}\ d\mu_{g}=-\epsilon^{4}\int_{{\mathbb{R}}^{n}}L_{0}VVdz+o(\epsilon^{4})$
Meanwhile, using the fact that $\Delta_{g}V-\Delta_{{\mathbb{R}}^{n}}V=o(\epsilon^{2})$ and rewriting the second term of (84) on ${\mathbb{R}}^{n}$:
@@
Hence, by adding (85) and (86), (84) turns into
@@
The first term, $\int_{{\mathbb{R}}^{n}}L_{0}V_{1}V_{1}dz$, can be computed as follows, using eq. (27) for $L_{0}V_{1}$ and eq. (29) for $V_{1}$,
$\epsilon^{4}\frac{1}{2}\int_{{\mathbb{R}}^{n}}L_{0}V_{1}V_{1}dz=-\epsilon^{4}\frac{1}{2}\frac{1}{9}\left(R_{kl}R_{ij}\int_{{\mathbb{R}}^{n}}\psi\frac{U^{\prime}}{|z|}z_{k}z_{l}z_{i}z_{j}dz\right)$
@@
since $\psi\frac{U^{\prime}}{|z|}$ is radial; with $c_{3}$ a dimensional constant, given by
(90) $c_{3}:=\frac{1}{54}\int_{{\mathbb{R}}^{n}}\psi(|z|)\frac{U^{\prime}(|z|)}{|z|}z_{1}^{4}dz,$
where $\psi=\psi(|z|)$ is the radial function discussed in Section 2, eqs. (28) and (29). Note that for (89), we have used the fact that for any radial function $\varphi(|z|)$ in ${\mathbb{R}}^{n}$,
$\int_{\mathbb{R}^{n}}\varphi(|z|)z_{i}z_{j}z_{k}z_{l}dz=(\delta_{ij}\delta_{kl}+\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk})\int_{\mathbb{R}^{n}}\varphi(|z|)z_{1}^{2}z_{2}^{2}$
$\frac{1}{3}(\delta_{ij}\delta_{kl}+\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk})\int_{\mathbb{R}^{n}}\varphi(|z|)z_{1}^{4}$
which follows from the fact that
$\int_{\mathbb{S}^{n-1}}z_{i}z_{j}z_{k}z_{l}=(\delta_{ij}\delta_{kl}+\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk})\int_{\mathbb{S}^{n-1}}z_{1}^{2}z_{2}^{2}$
$\frac{1}{3}(\delta_{ij}\delta_{kl}+\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk})\int_{\mathbb{S}^{n-1}}z_{1}^{4}$
In a similar way, for any radial function $\varphi(|z|)$ on ${\mathbb{R}}^{n}$,
(91) $\int_{{\mathbb{R}}^{n}}\varphi(|z|)z_{k}z_{l}dz=\delta_{kl}\int_{{\mathbb{R}}^{n}}\varphi(|z|)z_{1}^{2}dz$
We now compute the rest of the terms in (88). For the second term in (88), by direct computation, using eq. (27) for $L_{0}V_{1}$ and eq. (31) for $V_{2}$,
@@
$\frac{{\bf c}}{3}s_{g}^{2}\int_{{\mathbb{R}}^{n}}\left(\frac{1}{2}(U^{\prime}(z))^{2}-\frac{1}{2-p}\frac{U(z)U^{\prime}(z)}{|z|}\right)z_{1}^{2}dz$
by (91), since $\left(\frac{1}{2}(U^{\prime}(z))^{2}-\frac{1}{2-p}\frac{U(z)U^{\prime}(z)}{|z|}\right)$ is a radial function. On the other hand, by eq. (26) for $L_{0}V_{2}$ and eq. (29) for $V_{1}$,
$\int_{{\mathbb{R}}^{n}}L_{0}V_{2}V_{1}dz=-\frac{1}{3}{\bf c}s_{g}R_{kl}\int_{{\mathbb{R}}^{n}}U(z)\psi(|z|)z_{k}z_{l}dz$
$\frac{1}{3}{\bf c}s_{g}R_{kl}\delta_{kl}\int_{{\mathbb{R}}^{n}}U(z)\psi(|z|)z_{1}^{2}dz=-\frac{1}{3}{\bf c}s_{g}^{2}\int_{{\mathbb{R}}^{n}}U(z)\psi(|z|)z_{1}^{2}dz=2c_{4}s_{g}^{2}$
with
$c_{4}=-\frac{1}{6}{\bf c}\int_{{\mathbb{R}}^{n}}U(z)\psi(|z|)z_{1}^{2}dz$
since $U(z)\psi(|z|)$ a radial function. And by eq. (26) for $L_{0}V_{2}$ and eq. (31) for $V_{2}$:
@@
We then get, from these equations and (83), together with (87) and (88),
$K_{1}+K_{2}=\mathcal{A}(\xi)+o(\epsilon^{4})=-\epsilon^{4}\frac{1}{2}\int_{{\mathbb{R}}^{n}}L_{0}VVdz+o(\epsilon^{4})$
@@
(93) $c_{5}:=\frac{{\bf c}}{6}\int_{{\mathbb{R}}^{n}}\left(\frac{1}{2}(U^{\prime}(z))^{2}-\frac{1}{2-p}\frac{U(z)U^{\prime}(z)}{|z|}\right)z_{1}^{2}dz$
@@
On the other hand, $K_{3}$, $K_{4}$, $K_{5}$ and $K_{6}$ are $o(\epsilon^{4})$ by Lemmas 6.3, 6.4, 6.6 and 6.7, respectively. We conclude,
@@
where, $c_{6}=8c_{1}-120(n+2)c_{3}$, $c_{7}=-c_{3}-c_{4}-c_{5}-\frac{c_{2}\mathbf{c}}{12}+\frac{c_{1}}{24(n+2)}$ and $c_{8}=18c_{1}+30{\bf c}c_{2}(n+2)$.
We now are ready to prove Proposition 3.2.
@@
and the Proposition will follow from Lemma 4.1. We write $F(u)=\frac{1}{p}(u^{+})^{p}$ and $f(u)=(u^{+})^{p-1}$. First, since $\phi_{\epsilon,\bar{\xi}}\in K^{\perp}_{\epsilon,\bar{\xi}}$,
@@ 5. The Finite dimensional reduction
We follow [12, Proposition 3.1], [19, Lemma 6.1]. We proceed by contradiction. Suppose there is a sequence $\epsilon_{j}\rightarrow 0$, $\{\epsilon_{j}\}_{j\in{\mathbb{N}}}$ together with $\{\bar{\xi}_{j}\}_{j\in{\mathbb{N}}}\subset M^{K}$, such that the hypothesis of the Lemma, eq. (99) is satisfied and at the same time there is some $\{\phi_{j}\}_{j\in{\mathbb{N}}}\subset K^{\perp}_{\epsilon,\bar{\xi}}$, with